OpenStudy (anonymous):

An arc of a circle is of length 5p cm and the sector is bounds has an area of 20p cm^2 find the radius of the circle.

OpenStudy (sirm3d):

\[A_{sector}=(1/2)r^2 \theta=(1/2)r (r\theta),\quad s=r\theta\]